# One latent to fit them all: a unified representation of baryonic feedback on matter distribution

### v2

@@ 3.4 How latents affect the spectrum
In this section, we use the correlation patterns in Section 3.3 to interpret the latent effects on $T^{2}$ reconstruction (Section 3.2) and uncover their physical meaning.
Based on Section 3.3, “Latent 0” is positively correlated with the BH feedback via (20) BH Radiative Efficiency – (21) Quasar Threshold, while a large “Latent 1” suppresses the BH feedback through combined effects from (0) Wind Energy, (6) IMF Slope, (20) BH Radiative Efficiency, and (21) Quasar Threshold. To explore this, we examine the fraction of massive black holes ($M_{\mathrm{BH}}\geq 10^{8}M_{\odot}$, denoted as $f_{\mathrm{massive\ BH}}$), across different redshifts in the IllustrisTNG suite, finding consistent trends between BH mass and the latents across all redshifts, as shown in the upper panel of Figure 4.
@@
3.5 Stability of latent space dimension
@@
Even though our model with a two-dimensional latent space works quite well on the general datasets, we further test whether a higher-dimensional latent space would lead to better performance. We first performed a Principal Component Analysis (PCA) for the transfer functions. For all suites, the explained variance fractions for the first three components are $70\sim 80\%$, $10\sim 15\%$, and $5\%$, respectively, indicating that only the first two components are significant. As PCA is computed directly with $T^{2}$, it contains cosmology. Thus, the number of principal components should be no less than our latent dimensionality, supporting our 2D latent space.
To confirm that two latent dimensions suffice, we further train 3D models using the same strategy as in Section 2.3. However, the models on the Pareto front exhibit three types of misbehavior as shown in Figure 5: (1) incomplete reconstruction, in the top left panel, (2) an oddly behaving latent space (either overly dispersed or highly concentrated) in the top right panel, and (3) a failure to disentangle, which is shown in the bottom panel, where two latent dimensions are strongly correlated or there is one meaningless latent. The first two indicate overfitting, while the third suggests the model attempts to replicate an existing 2D latent direction rather than learning a new one. This demonstrates the model’s inability to form an independent third latent.
Combining these results with the physical interpretation of the 2D latent space in Section 3.4, we conclude that two latent dimensions are sufficient to capture baryonic physics across the four CAMELS suites.
4 Discussion
@@
At present, our architecture shows biases when SIMBA is included in the training set, likely due to its extreme feedback model, and our physical interpretation remains largely driven by IllustrisTNG. These limitations could be alleviated with the upcoming second-generation CAMELS simulations, which feature larger volumes and more sophisticated baryonic physics. A broader and more diverse dataset would enable the model to capture more complex and extreme feedback phenomena. In parallel, more advanced architectural extensions could be explored once supported by richer training data.
@@
Appendix A CAMELS simulations
As introduced in Section 2.1, we use four suites from CAMELS: IllustrisTNG SB28, Astrid SB7, SIMBA LH6, and Swift-EAGLE LH6. Here we summarized the basic properties of the four suites in Table 1.
@@ 22 Quasar Threshold Power 23–27 Cosmological Parameters
Appendix B Deduction of the ELBO of VAE
B.1 The family of VAE
@@
$\mathscr{L}_{\mathrm{VAE}}=\frac{1}{N}\sum_{n=1}^{N}(\mathbb{E}_{q(\bm{z}|\bm{x}_{n})}[\log p(\bm{x}_{n}|\bm{z})]-\operatorname{KL}(q(\bm{z}|\bm{x}_{n})||p(\bm{z}))).$ (B1)
@@
$\mathscr{L}_{\beta}=\frac{1}{N}\sum_{n=1}^{N}\left(\mathbb{E}_{q}[\log p(\bm{x}_{n}|\bm{z})]-\beta*\operatorname{KL}(q(\bm{z}|\bm{x}_{n})||p(\bm{z}))\right).$ (B2)
@@
$\frac{1}{N}\sum_{n=1}^{N}(\operatorname{KL}(q(\bm{z}|\bm{x}_{n})||p(\bm{z})))$ (B3)
$=$ $\mathbb{E}_{p(\bm{x}_{n})}[\operatorname{KL}(q(\bm{z}|\bm{x}_{n})||p(\bm{z}))]$ (B4)
$=$ $\mathbb{E}_{p(\bm{x}_{n})}[\mathbb{E}_{q(\bm{z}|\bm{x}_{n})}[\log(q(\bm{z}|\bm{x}_{n}))-\log(p(\bm{z}))+$ (B5)
$\log(q(\bm{z}))-\log(q(\bm{z}))+\log(\prod_{j}q(\bm{z}_{j}))-\log(\prod_{j}q(\bm{z}_{j}))]]$ (B6)
$=$ $\mathbb{E}_{q(\bm{z},\bm{x}_{n})}[\log(\frac{q(\bm{z}|x_{n})}{q(\bm{z})})]+\mathbb{E}_{q(\bm{z})}[\log(\frac{q(\bm{z})}{\prod_{j}q(z_{j})})]+\mathbb{E}_{q(\bm{z})}[\log(\frac{\prod_{j}q(z_{j})}{p(\bm{z})})]$ (B7)
$=$ $\mathbb{E}_{q(\bm{z},\bm{x}_{n})}[\log(\frac{q(\bm{z},\bm{x}_{n})}{q(\bm{z})p(\bm{x}_{n})})]+\mathbb{E}_{q(\bm{z})}[\log(\frac{q(\bm{z})}{\prod_{j}q(z_{j})})]+\sum_{j}\mathbb{E}_{q(z_{j})}[\log(\frac{q(z_{j})}{p(z_{j})})]$ (B8)
$=$ $\operatorname{KL}(q(\bm{z},\bm{x}_{n})||q(\bm{z})p(\bm{x}_{n}))+\operatorname{KL}(q(\bm{z})||\prod_{j}q(z_{j}))+\sum_{j}\operatorname{KL}(q(z_{j})||p(z_{j})).$ (B9)
These are the three KL term introduced in Section 2.2. As we would like to find an explainable latent representation for the baryonic feedback effects with the transfer function, we choose $\beta$-TCVAE to seek a set of disentangled latents that have desired information with feedback effects.
B.2 Reconstruction Loss
@@
$p(\bm{x}_{n}|\bm{z})=\mathcal{N}(\mu_{\theta}(\bm{z}),\sigma_{\theta}(\bm{z})^{2}).$ (B11)
@@
$\sigma_{\theta,i}^{2}=\mathrm{MSE}(\bm{x}^{\prime}_{i},\bm{x}_{i}),$ (B12)
@@
$\mathbb{E}_{q(\bm{z}|\bm{x}_{n})p(\bm{x}_{n})}[\log p(\bm{x}_{n}|\bm{z})]=-\sum_{i}\left(\log\sigma_{\theta,i}+\frac{1}{2\sigma_{\theta,i}^{2}}\,\mathrm{MSE}(\bm{x}^{\prime}_{i},\bm{x}_{i})\right)$ (B13)
@@
Combining the results from Equation B9 and Equation B13, the ELBO for $\beta$-TCVAE can be written as:
@@
$-\alpha\operatorname{KL}\bigl{(}q(\bm{z},\bm{x}_{n})\mathrel{\|}q(\bm{z})p(\bm{x}_{n})\bigr{)}$
@@
$+\gamma\sum_{j}\operatorname{KL}\bigl{(}q(z_{j})\mathrel{\|}p(z_{j})\bigr{)}.$ (B14)
Equation B14 generalizes VAE, which is a special case when $\alpha=\beta=\gamma=1$. And if $\alpha=\beta=\gamma>1$, it becomes $\beta$-VAE. Appendix C details the loss computation methods.
Appendix C Calculation of the Losses
C.1 Sampling in Minibatch
@@
$f(\bm{z},n*,\hat{B}_{M})=\frac{1}{N}q(\bm{z}|\bm{x}_{n*})+\frac{1}{M-1}\sum_{m=1}^{M-2}q(\bm{z}|\bm{x}_{m})+\frac{N-(M-1)}{N(M-1)}q(\bm{z}|\bm{x}_{M-1}),$ (C1)
@@
$\bar{f}(\bm{z},n*,\hat{B}_{M})=\frac{1}{N}q(\bm{z}|\bm{x}_{n*})+\frac{N-1}{N(M-1)}\sum_{m=1}^{M-1}q(\bm{z}|\bm{x}_{m}).$ (C3)
It can be easily seen that Equation C3 returns to the definition of $q(\bm{z})$ (i.e., $q(\bm{z})=\frac{1}{N}\sum_{n}q(\bm{z}|\bm{x}_{n})$) when $M=N$.
C.2 Calculation of KL terms
@@
$\mathscr{L}_{\mathrm{MI}}$ $=\mathbb{E}_{q(\bm{z},\bm{x}_{n})}[\log(\frac{q(\bm{z}|\bm{x})}{q(\bm{z})})]$ (C4)
@@
$\mathscr{L}_{\mathrm{TC}}$ $=\int q(\bm{z})\ln\left(\frac{q(\bm{z})}{\prod_{j}q(z_{j})}\right)\mathrm{d}z$ (C5)
@@
In practice, as we would like a set of disentangled parameters, $\mathscr{L}_{\mathrm{TC}}$ would be close to 0. In this case, as $q(\bm{z}_{n})$ and $\prod_{j}q(z_{n,j})$ are not perfectly normalized due to sampling, the difference in their normalization may result in a negative $\mathscr{L}_{\mathrm{TC}}$. However, adding the normalization factor into the estimator would overestimate the TC loss. As a result, we decided to keep the unbiased form of Equation C5 while keeping in mind that a small negative $\mathscr{L}_{\mathrm{TC}}$ might occur.
C.3 Cyclical Annealing
@@
$\mathscr{L}^{\prime}_{\beta-\mathrm{TC}}$ (C6)
@@
Appendix D Model Architecture
@@
The decoder takes the sampled latent variables together with the same five cosmological parameters as input. These inputs are first processed by another MLP, which produces an intermediate representation. This intermediate output is reshaped to match the number of $k$-bins in the power spectrum and then passed through a pointwise convolution layer to reconstruct $T^{2}$. To incorporate the variational decoder as mentioned in Appendix B.2, we also output the variance of the reconstructed $T^{2}$, $\sigma_{\theta,j}=\mathrm{MSE}(\bm{x}^{\prime}_{j},\bm{x}_{j})$. This approach allows the model to learn a more flexible and accurate representation of the data by considering the variance in the reconstruction process (O. Rybkin et al., 2021).
@@
Figure 6: An example of the training and validation loss curves. We use a learning rate scheduler with a patience of 1000 epochs and a 1% threshold, along with early stopping if no improvement is observed within 2000 epochs. Due to the cyclic annealing strategy described in Appendix C.3, the KL and reconstruction losses may exhibit oscillations during training, but they generally converge to stable values by the end.
@@
Appendix E Model Tuning
As noted in Section 2.3, we optimize 21 hyperparameters related to the model architecture, training procedure, and loss formulation using the Python package OPTUNA. A complete list of these hyperparameters is provided in Table 3. In addition to the optimized settings, we fix the batch size to match the size of the training (or validation) set to avoid bias in loss estimation that can arise from small batches.
@@ Appendix E Model Tuning
$1\leq A=\frac{\mathrm{Var}(\mu(\bm{x}))}{\overline{\sigma(\bm{x})^{2}}}\leq 10,$ (E1)
@@
Appendix F Results with SIMBA
@@
As discussed in Section 2.3, our general training dataset for the TEA model does not include SIMBA LH6 due to its strong baryonic feedback. In this section, we present results involving the SIMBA suite to further justify the construction of the “TEA” model.
We compare two models. The first, denoted as EAST, is trained on all four simulation suites (IllustrisTNG, Astrid, SIMBA, and EAGLE) using the same hyperparameter setup described in Appendix E. The second, called the “TEA tuned” model, starts from the EAST model and is further fine-tuned on the TEA dataset, which excludes SIMBA.
@@
This issue is further illustrated in Figure 9, which mirrors the latent variation analysis shown in the middle-left panel of Figure 1. The EAST model shows strong reconstruction bias at low $k$ in the Astrid suite, and the latent dimensions exert a large influence in that regime. After fine-tuning on the TEA dataset, the “TEA tuned” model significantly reduces this bias, particularly in the large-scale regime. Based on these results, we eventually decide to use TEA model for the main analysis shown in this paper.
Appendix G Reproducibility Test

### v3

@@ III.4 How latents affect the spectrum
In this section, we use the correlation patterns in Section III.3 to interpret the latent effects on $T^{2}$ reconstruction (Section III.2) and uncover their physical meaning.
Based on Section III.3, “Latent 0” is positively correlated with the BH feedback via (20) BH Radiative Efficiency – (21) Quasar Threshold, while a large “Latent 1” suppresses the BH feedback through combined effects from (0) Wind Energy, (6) IMF Slope, (20) BH Radiative Efficiency, and (21) Quasar Threshold. To explore this, we examine the fraction of massive black holes ($M_{\mathrm{BH}}\geq 10^{8}M_{\odot}$, denoted as $f_{\mathrm{massive\ BH}}$), across different redshifts in the IllustrisTNG suite, finding consistent trends between BH mass and the latents across all redshifts, as shown in the upper panel of Figure 4.
@@
III.5 Stability of latent space dimension
@@
Even though our model with a two-dimensional latent space works quite well on the general datasets, we further test whether a higher-dimensional latent space would lead to better performance. We first performed a principal component analysis (PCA) for the transfer functions. For all suites, the explained variance fractions for the first three components are $70\sim 80\%$, $10\sim 15\%$, and $5\%$, respectively, indicating that only the first two components are significant, which is consistent with early PCA results (H.-J. Huang et al., 2019). As PCA is computed directly with $T^{2}$, it contains cosmology. Thus, the number of principal components should be no less than our latent dimensionality, supporting our 2D latent space. As for the reconstruction comparison between PCA and VAE, more details can be found in Appendix I.
To confirm that two latent dimensions suffice, we further train 3D models using the same strategy as in Section II.3. However, the models on the Pareto front exhibit three types of misbehavior as shown in Figure 5: (1) incomplete reconstruction, in the top left panel, (2) an oddly behaving latent space (either overly dispersed or highly concentrated) in the top right panel, and (3) a failure to disentangle, which is shown in the bottom panel, where two latent dimensions are strongly correlated or there is one meaningless latent. The first two indicate overfitting, while the third suggests the model attempts to replicate an existing 2D latent direction rather than learning a new one. This demonstrates the model’s inability to form an independent third latent.
Combining these results with the physical interpretation of the 2D latent space in Section III.4, we conclude that two latent dimensions are sufficient to capture baryonic physics across the four CAMELS suites.
IV Discussion
@@
At present, our architecture shows biases when SIMBA is included in the training set, likely due to its extreme feedback model, and our physical interpretation remains largely driven by IllustrisTNG. These limitations could be alleviated with the upcoming second-generation CAMELS simulations, which feature larger volumes and more sophisticated baryonic physics. However, as shown in Figure 2, the TEA model can still reach a decent reconstruction on SIMBA within the range expected from cosmic variance. On the other hand, though we already include over $5,000$ simulations from 4 different suites, a broader and more diverse dataset would enable the model to capture more complex and extreme feedback phenomena. In parallel, more advanced architectural extensions could be explored once supported by richer training data.
@@
Appendix A Comparison between Methods
Table 1: Comparison across five representative methods: baryonic correction model (BCM), symbolic regression (SR), physically motivated parameterizations (PMP), principal component analysis (PCA), and our $\beta$-TCVAE method. $\checkmark$ = yes and $\times$ = no. D means data-driven, which typically uncover more parsimonious representations, and M stands for hand-crafted, that could result in large parameterizations.
Criterion BCM SR PMP PCA VAE (this work)
Driven by Data/Modeler M D M D D
Subgrid Model Independence $\times$
Cosmology Independence $\times$ $\times$
Redshift Independence $\times$
Interpretability $\times$ $\times$
Disentangled Representation $\times$ $\times$ $\times$ $\times$
As discussed in section I, numerous studies have investigated baryonic feedback effects using the transfer function. Here, we summarize the advantages and limitations of five methods: baryonic correction models (BCM), symbolic regression (SR), physically motivated parameterizations (PMP), principal component analysis (PCA), and our $\beta$ total correlation variational autoencoders ($\beta$-TCVAE) in Table 1.
Appendix B CAMELS simulations
As introduced in Section II.1, we use four suites from CAMELS: IllustrisTNG SB28, Astrid SB7, SIMBA LH6, and Swift-EAGLE LH6. Here we summarized the basic properties of the four suites in Table 2.
@@ 22 Quasar Threshold Power 23–27 Cosmological Parameters
Appendix C Deduction of the ELBO of VAE
C.1 The family of VAE
@@
$\mathscr{L}_{\mathrm{VAE}}=\frac{1}{N}\sum_{n=1}^{N}(\mathbb{E}_{q(\bm{z}|\bm{x}_{n})}[\log p(\bm{x}_{n}|\bm{z})]-\operatorname{KL}(q(\bm{z}|\bm{x}_{n})||p(\bm{z}))).$ (C1)
@@
$\mathscr{L}_{\beta}=\frac{1}{N}\sum_{n=1}^{N}\left(\mathbb{E}_{q}[\log p(\bm{x}_{n}|\bm{z})]-\beta*\operatorname{KL}(q(\bm{z}|\bm{x}_{n})||p(\bm{z}))\right).$ (C2)
@@
$\frac{1}{N}\sum_{n=1}^{N}(\operatorname{KL}(q(\bm{z}|\bm{x}_{n})||p(\bm{z})))$ (C3)
$=$ $\mathbb{E}_{p(\bm{x}_{n})}[\operatorname{KL}(q(\bm{z}|\bm{x}_{n})||p(\bm{z}))]$ (C4)
$=$ $\mathbb{E}_{p(\bm{x}_{n})}[\mathbb{E}_{q(\bm{z}|\bm{x}_{n})}[\log(q(\bm{z}|\bm{x}_{n}))-\log(p(\bm{z}))+$ (C5)
$\log(q(\bm{z}))-\log(q(\bm{z}))+\log(\prod_{j}q(\bm{z}_{j}))-\log(\prod_{j}q(\bm{z}_{j}))]]$ (C6)
$=$ $\mathbb{E}_{q(\bm{z},\bm{x}_{n})}[\log(\frac{q(\bm{z}|x_{n})}{q(\bm{z})})]+\mathbb{E}_{q(\bm{z})}[\log(\frac{q(\bm{z})}{\prod_{j}q(z_{j})})]+\mathbb{E}_{q(\bm{z})}[\log(\frac{\prod_{j}q(z_{j})}{p(\bm{z})})]$ (C7)
$=$ $\mathbb{E}_{q(\bm{z},\bm{x}_{n})}[\log(\frac{q(\bm{z},\bm{x}_{n})}{q(\bm{z})p(\bm{x}_{n})})]+\mathbb{E}_{q(\bm{z})}[\log(\frac{q(\bm{z})}{\prod_{j}q(z_{j})})]+\sum_{j}\mathbb{E}_{q(z_{j})}[\log(\frac{q(z_{j})}{p(z_{j})})]$ (C8)
$=$ $\operatorname{KL}(q(\bm{z},\bm{x}_{n})||q(\bm{z})p(\bm{x}_{n}))+\operatorname{KL}(q(\bm{z})||\prod_{j}q(z_{j}))+\sum_{j}\operatorname{KL}(q(z_{j})||p(z_{j})).$ (C9)
These are the three KL term introduced in Section II.2. As we would like to find an explainable latent representation for the baryonic feedback effects with the transfer function, we choose $\beta$-TCVAE to seek a set of disentangled latents that have desired information with feedback effects.
C.2 Reconstruction Loss
@@
$p(\bm{x}_{n}|\bm{z})=\mathcal{N}(\mu_{\theta}(\bm{z}),\sigma_{\theta}(\bm{z})^{2}).$ (C11)
@@
$\sigma_{\theta,i}^{2}=\mathrm{MSE}(\bm{x}^{\prime}_{i},\bm{x}_{i}),$ (C12)
@@
$\mathbb{E}_{q(\bm{z}|\bm{x}_{n})p(\bm{x}_{n})}[\log p(\bm{x}_{n}|\bm{z})]=-\sum_{i}\left(\log\sigma_{\theta,i}+\frac{1}{2\sigma_{\theta,i}^{2}}\,\mathrm{MSE}(\bm{x}^{\prime}_{i},\bm{x}_{i})\right)$ (C13)
@@
Combining the results from Equation C9 and Equation C13, the ELBO for $\beta$-TCVAE can be written as:
@@
$-[\alpha\operatorname{KL}\bigl(q(\bm{z},\bm{x}_{n})\mathrel{\|}q(\bm{z})p(\bm{x}_{n})\bigr)$
@@
$+\gamma\sum_{j}\operatorname{KL}\bigl(q(z_{j})\mathrel{\|}p(z_{j})\bigr)].$ (C14)
Equation C14 generalizes VAE, which is a special case when $\alpha=\beta=\gamma=1$. And if $\alpha=\beta=\gamma>1$, it becomes $\beta$-VAE. Appendix D details the loss computation methods.
Appendix D Calculation of the Losses
D.1 Sampling in Minibatch
@@
$f(\bm{z},n*,\hat{B}_{M})=\frac{1}{N}q(\bm{z}|\bm{x}_{n*})+\frac{1}{M-1}\sum_{m=1}^{M-2}q(\bm{z}|\bm{x}_{m})+\frac{N-(M-1)}{N(M-1)}q(\bm{z}|\bm{x}_{M-1}),$ (D1)
@@
$\bar{f}(\bm{z},n*,\hat{B}_{M})=\frac{1}{N}q(\bm{z}|\bm{x}_{n*})+\frac{N-1}{N(M-1)}\sum_{m=1}^{M-1}q(\bm{z}|\bm{x}_{m}).$ (D3)
It can be easily seen that Equation D3 returns to the definition of $q(\bm{z})$ (i.e., $q(\bm{z})=\frac{1}{N}\sum_{n}q(\bm{z}|\bm{x}_{n})$) when $M=N$.
D.2 Calculation of KL terms
@@
$\mathscr{L}_{\mathrm{MI}}$ $=\mathbb{E}_{q(\bm{z},\bm{x}_{n})}[\log(\frac{q(\bm{z}|\bm{x})}{q(\bm{z})})]$ (D4)
@@
$\mathscr{L}_{\mathrm{TC}}$ $=\int q(\bm{z})\ln\left(\frac{q(\bm{z})}{\prod_{j}q(z_{j})}\right)\mathrm{d}z$ (D5)
@@
In practice, as we would like a set of disentangled parameters, $\mathscr{L}_{\mathrm{TC}}$ would be close to 0. In this case, as $q(\bm{z}_{n})$ and $\prod_{j}q(z_{n,j})$ are not perfectly normalized due to sampling, the difference in their normalization may result in a negative $\mathscr{L}_{\mathrm{TC}}$. However, adding the normalization factor into the estimator would overestimate the TC loss. As a result, we decided to keep the unbiased form of Equation D5 while keeping in mind that a small negative $\mathscr{L}_{\mathrm{TC}}$ might occur.
D.3 Cyclical Annealing
@@
$\mathscr{L}^{\prime}_{\beta-\mathrm{TC}}$ (D6)
@@
Appendix E Model Architecture
@@
The decoder takes the sampled latent variables together with the same five cosmological parameters as input. These inputs are first processed by another MLP, which produces an intermediate representation. This intermediate output is reshaped to match the number of $k$-bins in the power spectrum and then passed through a pointwise convolution layer to reconstruct $T^{2}$. To incorporate the variational decoder as mentioned in Appendix C.2, we also output the variance of the reconstructed $T^{2}$, $\sigma_{\theta,j}=\mathrm{MSE}(\bm{x}^{\prime}_{j},\bm{x}_{j})$. This approach allows the model to learn a more flexible and accurate representation of the data by considering the variance in the reconstruction process (O. Rybkin et al., 2021).
@@
Figure 6: An example of the training and validation loss curves. We use a learning rate scheduler with a patience of 1000 epochs and a 1% threshold, along with early stopping if no improvement is observed within 2000 epochs. Due to the cyclic annealing strategy described in Appendix D.3, the KL and reconstruction losses may exhibit oscillations during training, but they generally converge to stable values by the end.
@@
Appendix F Model Tuning
As noted in Section II.3, we optimize 21 hyperparameters related to the model architecture, training procedure, and loss formulation using the Python package OPTUNA. A complete list of these hyperparameters is provided in Table 4. In addition to the optimized settings, we fix the batch size to match the size of the training (or validation) set to avoid bias in loss estimation that can arise from small batches.
@@ Appendix F Model Tuning
$1\leq A=\frac{\mathrm{Var}(\mu(\bm{x}))}{\overline{\sigma(\bm{x})^{2}}}\leq 10,$ (F1)
@@
Appendix G Results with SIMBA
@@
Figure 10: Range of $T^{2}$ for different suites. The blue region stands for the range between upper and lower limit in each $k$ bin. The red region represents the $16-84\%$ range in each $k$ bin. Comparing to the other suites, SIMBA both significantly deviate from 1 and have much larger variance on the largest scales.
As discussed in Section II.3, our general training dataset for the TEA model does not include SIMBA LH6 due to its strong baryonic feedback. In this section, we present results involving the SIMBA suite to further justify the construction of the “TEA” model.
We compare two models. The first, denoted as EAST, is trained on all four simulation suites (IllustrisTNG, Astrid, SIMBA, and EAGLE) using the same hyperparameter setup described in Appendix F. The second, called the “TEA tuned” model, starts from the EAST model and is further fine-tuned on the TEA dataset, which excludes SIMBA.
@@
This issue is further illustrated in Figure 9, which mirrors the latent variation analysis shown in the middle-left panel of Figure 1. The EAST model shows strong reconstruction bias at low $k$ in the Astrid suite, and the latent dimensions exert a large influence in that regime. After fine-tuning on the TEA dataset, the “TEA tuned” model significantly reduces this bias, particularly in the large-scale regime. A reason for this is that, SIMBA not only have a stronger feedback than the other suite on the largest scales, but also have larger variance due to the wide range of CAMELS, as shown in Figure 10. Based on these results, we eventually decide to use TEA model for the main analysis shown in this paper.
Appendix H Reproducibility Test